# QPET: A Versatile and Portable Quantity-of-Interest-Preservation Framework for Error-Bounded Lossy Compression

### v3

@@ 4. QPET Design Overview
In this section, we provide an overview of the proposed QPET (QoI-oriented Point-wise Error-bound Tuning) framework, describe the underlying QoI-preserving error-bounded lossy compression algorithm, and feature its integration into several existing error-bounded lossy compressors.
4.1. QPET composition
In Figure 1, we demonstrate the composition of the QPET framework that provides QoI-preservation functionality, presenting how it integrates and interacts with general modular scientific error-bounded lossy compression pipelines such as SZ (Liang et al., 2018b; Zhao et al., 2021; Liang et al., 2023; Liu et al., 2022, 2024) and SPERR (Li et al., 2023). Compared to existing QoI-preserving work (Jiao et al., 2022), QPET introduces several brand-new or improved strategies to cover a wide range of QoI constraints and deliver much higher compression ratios. The components of QPET are:
@@ 4.1. QPET composition
  • Error Bound Compressor: This module compresses and stores the point-wise error bounds (this is an optional module because some compression pipelines cannot support point-wise error bounds). QPET follows a lossy scheme described in (Liang et al., 2020; Jiao et al., 2022).
@@ 4.2. QPET-integrated Compression Algorithm
4.3. QPET Integration
Because the design of QPET is only based on error-bound estimation/tuning and post-hoc data validation/correction, it does not rely on certain data compression techniques and implementation details. As such, QPET has high portability and adaptability to many existing error-bounded lossy compression frameworks. In this work, we have integrated QPET into three state-of-the-art scientific error-bounded lossy compressors with two archetypes: SZ3 (Zhao et al., 2021; Liang et al., 2023), HPEZ (Liu et al., 2024), and SPERR (Li et al., 2023). SZ3 and HPEZ are interpolation-prediction-based error-bounded lossy compressors, and SPERR is a wavelet-transform-based error-bounded lossy compressor. In the integration of QPET to each data compressor, the error-bound estimation and tuning modules (according to the QoI function and error threshold) are added before the data compression process, and the data validation/correction module serves as a data post-processing step in the compression pipeline.
5. QPET Design Components
In this section, we present the design components of QPET introduced in Section 4. First, we discuss the core of QPET: providing optimized point-wise error bounds for the QoI-preserving lossy compression. Then, we explain how to auto-tune the global data error bound according to the computed point-wise error bounds to optimize the compression ratio. Last, we specify how to correct the outlier data values to strictly respect the QoI error threshold.
@@ 5.1. Computing Point-wise Error Bounds with Univariate QoI
Table 1. Notations
@@
Given second-order differentiable QoI function $f$, a global error bound $\epsilon_{g}$, and a QoI error threshold $t$, for each data point $x_{i}$, if $f^{\prime\prime}(x_{i})\neq 0$, an estimation for the best-fit data error bound $\epsilon_{x_{i}}$ of data point $x_{0}$ to fit its QoI error threshold will be $\min{\left(\epsilon_{g},\frac{\sqrt{|a|^{2}+2|b|t}-|a|}{|b|}\right)}$, in which $a=f^{\prime}(x_{i})$ and $b=f^{\prime\prime}(x_{i})$.
@@
Using $f_{2}(x)=x_{0}+a(x-x_{0})+\frac{b}{2}(x-x_{0})^{2}$ as an approximation of $f$, we first find the maximum $\epsilon$ so that $|x_{1}-x_{0}|\leq\epsilon$implies $|f(x_{1})-f(x_{0})|\approx|f_{2}(x_{1})-f_{2}(x_{0})|\leq|a(x_{1}-x_{0})+\frac{b}{2}(x_{1}-x_{0})^{2}|\leq t$. This is equivalent to maximizing $\epsilon$ which satisfies that $[-\epsilon,\epsilon]\subseteq A$, in which $A$ is the solution set of $|ax+\frac{b}{2}x^{2}|\leq t$.
@@
Theorem 5.1 and Theorem 5.2 cover all the cases for best-fit point-wise error-bound estimation. In Figure 2, based on different pointwise QoIs and QoI error tolerances, we plot the point-wise data error bounds for different data values from both analytical solutions and our proposed estimation method. For various function formats (Figure 2 (a-c)), QPET acquires accurate error bound estimations compared to the analytical solutions, even when $|f^{\prime\prime\prime}(x)|$ has large values (see Figure 2 (c) where $f^{\prime\prime\prime}(x)=-100\sin{10x}$). In Figure 2 (d), we show an example of $f(x)=x^{3}$, which has close-to-zero first- and second-order derivatives when $x$ is close to zero. While QPET may produce over-estimated error bounds in these cases, those over-estimated error bounds will be cropped by either the pre-given data error bound or the auto-tuned global error bound (to be detailed in Section 5.3) and thus do no harm to the compression process (notice that in Figure 2 we set $\epsilon_{g}$ as +$\infty$+ ∞ to focus on the estimation quality from QoI error bounds). As long as QPET presents good estimations for small error bounds (it actually does), it will be optimized for the compression ratio.
@@ 5.2. Computing Point-wise Error Bounds with Multivariate QoI
In other words, compressing each $x_{i}$ with an error bound $\epsilon_{i}$ will lead the QoI error to be constrained. In the next, we will first discuss optimizing $\left\{\epsilon_{i}\right\}$ when $F\left(x_{1},\dots,x_{n}\right)$ is the linear combination of univariate terms, i.e., the variables are separated. Then, we will show how to extend the methodology to generalized cases.
@@ 5.2.1. Multivariate QoI in variable-separated format
Then we can optimize the point-wise data error bound $\left\{\epsilon_{i}\right\}$ for set $\left\{x_{i}\right\}$ from those $\left\{t_{i}\right\}$ with Algorithm 2, as in this case we are just bounding point-wise QoI $f$ error within $t_{i}$. Therefore, by solving Eq. 5 we can find a way to optimize $\left\{\epsilon_{i}\right\}$. Regarding the optimization target of Eq. 5, intuitively, we can maximize $\sum_{i}t_{i}$. However, this optimization target sometimes leads to skewed distributions of $t_{i}$ and $\epsilon_{i}$ values over different $i$. Preliminary works (Jiao et al., 2022) and our experiments showed that point-wise error bounds with skewed distribution would cause the compression ratio to decrease significantly. Therefore, we select to maximize $\min t_{i}$, equivalent to ensure $t_{i}=t$ for each i. According to our preliminary experiments, this scheme is better than other intuitive strategies, such as aligning $\alpha_{i}t_{i}$. Moreover, if we denote $D_{i}$ as $f\left(x_{i}^{\prime}\right)-f\left(x_{i}\right)$, we will notice that $F\left(x_{1}^{\prime},\dots,x_{n}^{\prime}\right)-F\left(x_{1},\dots,x_{n}\right)=\sum_{i}{\alpha_{i}D_{i}}$. To this end, we can convert Eq. 5 to a more concise format:
@@
Admittedly, Theorem 5.4 needs certain preliminary conditions of $\left\{D_{i}\right\}$, which are not always true. For example, the variance of compression errors is dependent on several other factors such as data characteristics and input error bound, and $\left\{D_{i}\right\}$ are not theoretically independent of each other. Nevertheless, existing research (Zhao et al., 2021; Liu et al., 2022, 2024) showed that, in practical cases, distributions of $\left\{D_{i}\right\}$ are not severely distorted from the assumptions, exhibiting distributions that are close to uniform or Gaussian, and the autocorrelation of decompression errors drop rapidly as the compression accuracy increases (Liu et al., 2022). Moreover, to handle the unclear characteristics of $\left\{D_{i}\right\}$, Theorem 5.4 can still acquire adequate QoI error threshold estimations by selecting a conservative value of $c$. To optimize the compression ratio, Theorem 5.4 does not bring a strict guarantee for bounding all QoI errors (this is the same for Eq. 8). In Section 5.4, we will discuss how QPET strictly guarantees the QoI error threshold and can optimize the compression ratio through more aggressive data error-bound settings.
@@ 5.2.2. Multivariate QoI in non-variable-separated format
Beyond the variable-separated format described in Eq. 4, regarding non-variable-separated multivariate QoI function $F\left(x_{1},\dots,x_{n}\right)$, to separate the variables so that techniques described in Section 5.2.1 can be applied, we use its approximation for computing point-wise error bounds, leveraging the first-order differentials of $F$ (using higher-order differentials will bring cross partial derivatives, failing in separate the variables). Specifically, we have:
@@
Eq. 8 falls into the format of Eq. 4, with $C=F\left(x_{1},\dots,x_{n}\right)$, $\alpha_{i}=\frac{\partial F}{\partial x_{i}}\left(x_{i}\right)$, and $f\left(x\right)=x-x_{i}$ (equivalent to $f(x)=x$). Therefore, we can use the Eq. 8 and Theorem 5.3/5.4 to determine the point-wise error bounds from $F$. It is worth noticing that, in this case, for different $\left\{x_{i}\right\}$, we will have different $t$ from Theorem 5.3 and  5.4 as there are different $\alpha_{i}$ values from the partial derivatives.
In Algorithm 3, we combine the computation of point-wise data error bounds from multivariate QoI in either linear or non-linear format. When dealing with the linear format, the computation can often get simplified. For example, for linear multivariate QoIs, the $t$ in Line 14 can be shared over data regions.
@@ 5.4. QoI error validation and correction
As previously described in Section 3.2, the design purpose of QPET is to optimize the compression ratio when strictly constraining the QoI error threshold. To fulfill this target, the point-wise and global error-bound calculations demonstrated in Section 5.1, Section 5.2, and Section 5.3 do not conservatively make the QoI value strictly-preserved on each data element, but require a subsequent process to correct and store the out-of-bound data points (i.e., outliers). By selecting proper error-bound values, QPET can significantly reduce the average storage cost of each compressed data value, meanwhile introducing negligible overheads for correcting outliers as they are only a tiny portion ($<1\%$) of the whole input data. Specifically, QPET detects and corrects the outliers in the following scheme: Given the decompressed data (which are naturally acquired after the compression process), QPET computes the QoI values on it and compares them to the original QoI values. For out-of-bound univariate QoI errors, QPET losslessly encodes and stores all the corresponding original data values. For each out-of-bound multivariate QoI error, there are several related data values. To minimize the number of data values that are losslessly stored, QPET stores those values one by one and updates the QoI result after each value update until the QoI error gets bounded.
@@ 6.1.1. Experimental environment and datasets
We perform the evaluations on 6 real-world scientific datasets from diverse domains (details in Table 3). Experiments are operated on Purdue Anvil computing cluster (Song et al., 2022) (each node is equipped with two 64-core AMD EPYC 7763 CPUs and 512GB DDR4 memory).
@@ 6.1.2. Baselines
Besides the QPET-integrated compressors, we included several existing solutions for QoI-preserving scientific lossy compression in the evaluations, which are in 2 categories: 1) Parameter-search-based solutions: general-purpose compressors (SZ3/HPEZ/SPERR) cannot directly bound specified QoI errors, so we apply parameter-search methods on top of them to figure out the best-fit (yielding highest compression ratio) data error bound for preserving the QoI with given accuracy. The parameter-search methods, including binary-search-based and FraZ (Underwood et al., 2020), are from OptZConfig (Underwood et al., 2022), which is the state-of-the-art parameter-search toolkit for scientific lossy compression. When applying, we slightly revised those methods to get them better adapted and accelerated in the QoI-preserving tasks. Those baselines are named SZ3/HPEZ/SPERR-OptZ-R (R is short for revised), and they represent the best (fastest) results from both parameter-search methods on every base compressor. In the parameter-search process, multiple iterations of data compression and QoI validation are performed to fit the QoI error bound, so they are typically quite slower than the original general-purpose compressors. (2) QoI-preserving compressors: (Jiao et al., 2022) provided the SZ3-based QoI-preserving compressor QoI-SZ3, and we further ported its QoI-preserving features to HPEZ, creating QoI-HPEZ. Moreover, we evaluated MGARD-QoI (Ainsworth et al., 2019). Those QoI-preserving compressors have limited support for diverse QoI formats. QoI-SZ3/HPEZ only supports square and logarithm (also their block-averages), and MGARD-QoI only supports linear QoIs. So, we only tested them on the QoIs supported. Other existing QoI-integrated compressors are designed for different tasks (e.g., cpSZ (Liang et al., 2018a) only works for critical points in vector field data), so they are not included in our evaluation baselines.
@@ 6.1.3. QoI functions, experimental configurations, and evaluation metrics
Table 4 shows the QoI functions in the evaluation tasks. Among them, there are three different categories: point-wise, regional, and vector. They have diverse mathematical formats, and for many among them (such as $\tanh{x}$, $\frac{1}{n}\sum{x^{3}}$, and vector QoIs), QPET is the first framework that supports compression with preservation of those QoIs. The selection of QoI functions in our evaluation is based on existing investigations and analysis (Jiao et al., 2022; Su et al., 2022; Wu et al., 2024) of QoIs in practical scientific data analysis tasks, such as physical transform (kinetic energy as velocity’s square), and clustering ($\sqrt{x^{2}+y^{2}+z^{2}}$ is the distance from origin when $x$, $y$, and $z$ are coordinates).
@@
For the OptZ-R parameter search, we set an early termination that triggers when a maximum QoI error between 90% and 100% of the required threshold is found because it presents a near-optimal compression ratio with reasonable search time. Regarding the compression configurations, we apply the default optimization level and compression-ratio-preferred mode for HPEZ. Regarding QPET parameters in Algorithms 3, we set $c=3$, $\beta=0.999$ for SPERR, $c=2$, $\beta=0.999$ for HPEZ, and $c=2$, $\beta=0.99999$ for SZ3. On SZ3 and HPEZ, the autocorrelation of decompression errors are relatively high when the error bound is large (Liu et al., 2022), so we linearly dynamically decrease $c$ when $\tau$ increases over $10^{-3}$, eventually to 1.0 when $\tau$ becomes $10^{-2}$.

### v4

@@ 4. QPET Design Overview
In this section, we provide an overview of the proposed QPET (QoI-oriented Point-wise Error-bound Tuning) framework, describing its modules, underlying algorithm, and its integration into several existing error-bounded lossy compression pipelines.
4.1. QPET composition and algorithm
In Figure 1, we demonstrate the composition of the QPET framework that provides QoI-preservation functionality, presenting how it integrates and interacts with general modular scientific error-bounded lossy compression pipelines. Compared to existing QoI-preserving work (Jiao et al., 2022), QPET introduces several brand-new or improved strategies to cover a wide range of QoI constraints and deliver high compression ratios. The components of QPET are:
@@ 4.1. QPET composition and algorithm
  • Error Bound Compressor: This module compresses and stores the point-wise error bounds (will be skipped if the base compression pipeline cannot support point-wise error bounds) with improvements from (Liang et al., 2020; Jiao et al., 2022).
@@
4.2. Integrating QPET into error-bounded compression pipelines
Because the design of QPET is only based on error-bound estimation/tuning and post-hoc data validation/correction, it does not rely on certain data compression techniques and implementation details. As such, QPET has high portability and adaptability to many existing error-bounded lossy compression pipelines. In this work, we have integrated QPET into three state-of-the-art scientific error-bounded lossy compressors with two archetypes: SZ3 (Zhao et al., 2021; Liang et al., 2023), HPEZ (Liu et al., 2024), and SPERR (Li et al., 2023). SZ3 and HPEZ are interpolation-prediction-based error-bounded lossy compressors, and SPERR is a wavelet-transform-based error-bounded lossy compressor. In the integration of QPET to each data compressor, the error-bound estimation and tuning modules (according to the QoI function and error threshold) are added before the data compression process, and the data validation/correction module serves as a data post-processing step in the compression pipeline.
5. QPET Design Details
In this section, we present the design details of QPET introduced in Section 4. First, we discuss the core of QPET: providing optimized point-wise error bounds for the QoI-preserving lossy compression. Then, we explain how to auto-tune the global data error bound according to the computed point-wise error bounds to optimize the compression ratio. Last, we specify how to correct the outlier data values to strictly respect the QoI error threshold.
@@ 5.1. Computing Point-wise Error Bounds with Univariate QoI
Table 1. Mathematical notations for QoI-preserving
@@
Given a second-order differentiable QoI function $f$, a global error bound $\epsilon_{g}$, and a QoI error threshold $t$, for each data point $x_{i}$, if $f^{\prime\prime}(x_{i})\neq 0$, an estimation for the best-fit data error bound $\epsilon_{x_{i}}$ to fit its QoI error threshold will be $\min{\left(\epsilon_{g},\frac{\sqrt{|a|^{2}+2|b|t}-|a|}{|b|}\right)}$, in which $a=f^{\prime}(x_{i})$ and $b=f^{\prime\prime}(x_{i})$.
@@
Using $f_{2}(x)=x_{i}+a(x-x_{i})+\frac{b}{2}(x-x_{i})^{2}$ as an approximation of $f$, we first find the maximum $\epsilon$ so that $|x_{i}^{\prime}-x_{i}|\leq\epsilon$implies $|f(x_{i}^{\prime})-f(x_{i})|\approx|f_{2}(x_{i}^{\prime})-f_{2}(x_{i})|\leq|a(x_{i}^{\prime}-x_{i})+\frac{b}{2}(x_{i}^{\prime}-x_{i})^{2}|\leq t$. This is equivalent to maximizing $\epsilon$ which satisfies that $[-\epsilon,\epsilon]\subseteq A$, in which $A$ is the solution set of $|ax+\frac{b}{2}x^{2}|\leq t$.
@@
Theorem 5.1 and Theorem 5.2 cover all the cases for point-wise error-bound estimation. In Figure 2, based on different pointwise QoIs and error tolerances, we plot the point-wise data error bounds for different data values from both analytical solutions and our proposed estimation method. For various function formats (Figure 2 (a-c)), QPET acquires accurate error bound estimations compared to the analytical solutions, even when $|f^{\prime\prime\prime}(x)|$ has large values (see Figure 2 (c) where $f^{\prime\prime\prime}(x)=-1000\cos{10x}$). In Figure 2 (d), we show an example of $f(x)=x^{3}$, which has close-to-zero first- and second-order derivatives when $x$ is close to zero. While QPET may produce over-estimated error bounds in this case, the over-estimated error bounds will be cropped by a pre-given or auto-tuned global error bound (to be detailed in Section 5.3) and thus not harm the compression process (notice that in Figure 2 we set $\epsilon_{g}$ as +$\infty$+ ∞ to focus on the estimation quality from QoI error bounds). As long as QPET presents good estimations for small error bounds (it does), it will be optimized for the compression ratio.
@@ 5.2. Computing Point-wise Error Bounds with Multivariate QoI
In other words, compressing each $x_{i}$ with an error bound $\epsilon_{i}$ will lead the QoI error to be constrained. In the following, we first discuss optimizing $\left\{\epsilon_{i}\right\}$ when $F\left(x_{1},\dots,x_{n}\right)$ is the linear combination of univariate terms, i.e., the variables are separated. Then, we will demonstrate how to extend the methodology to more general cases.
@@ 5.2.1. Multivariate QoI in variable-separated format
Then we can optimize the point-wise data error bound $\left\{\epsilon_{i}\right\}$ for set $\left\{x_{i}\right\}$ from those $\left\{t_{i}\right\}$ with Algorithm 2, as in this case we are just bounding point-wise QoI $f$ error within $t_{i}$. Therefore, by solving Eq. 5 we can find a way to optimize $\left\{\epsilon_{i}\right\}$. Regarding the optimization target of Eq. 5, intuitively, we can maximize $\sum_{i}t_{i}$. However, this optimization target sometimes leads to skewed distributions of $t_{i}$ and $\epsilon_{i}$ values over different $i$. Preliminary works (Jiao et al., 2022) and our experiments have shown that point-wise error bounds with skewed distributions would cause the compression ratio to decrease significantly. Therefore, we select to maximize $\min t_{i}$, equivalent to ensure $t_{i}=t$ for each i. According to our preliminary experiments, this scheme outperforms other intuitive strategies, such as aligning $\alpha_{i}t_{i}$. Moreover, if we denote $D_{i}$ as $f\left(x_{i}^{\prime}\right)-f\left(x_{i}\right)$, we will notice that $F\left(x_{1}^{\prime},\dots,x_{n}^{\prime}\right)-F\left(x_{1},\dots,x_{n}\right)=\sum_{i}{\alpha_{i}D_{i}}$. To this end, we can convert Eq. 5 to a more concise format:
@@
Admittedly, Theorem 5.4 needs certain preliminary conditions of $\left\{D_{i}\right\}$, which are not always true. The variance of compression errors is dependent on diverse factors, such as data characteristics and input error bound, and $\left\{D_{i}\right\}$ are not theoretically independent of each other (such as in interpolation-based data compression). Nevertheless, existing research (Zhao et al., 2021; Liu et al., 2022, 2024) showed that, in practical cases, distributions of $\left\{D_{i}\right\}$ are not severely distorted from the assumptions, exhibiting distributions that are close to uniform or Gaussian, and the autocorrelation of decompression errors drop rapidly as the compression accuracy increases (Liu et al., 2022). Moreover, to handle the unclear characteristics of $\left\{D_{i}\right\}$, Theorem 5.4 can still acquire adequate QoI error threshold estimations by selecting a conservative value of $c$. To optimize the compression ratio, Theorem 5.4 does not bring a strict guarantee for bounding all QoI errors. In Section 5.4, we will discuss how QPET strictly guarantees the QoI error threshold and can optimize the compression ratio through more aggressive data error-bound settings.
@@ 5.2.2. Multivariate QoI in non-variable-separated format
Beyond the variable-separated format described in Eq. 4, regarding a non-variable-separated multivariate QoI function $F\left(x_{1},\dots,x_{n}\right)$, to separate the variables so that techniques described in Section 5.2.1 can be applied, we use its approximation for computing point-wise error bounds, leveraging the first-order differentials of $F$ (using higher-order differentials will bring cross partial derivatives, failing to separate the variables). Specifically, we have:
@@
Eq. 8 falls into the format of Eq. 4, with $C=F\left(x_{1},\dots,x_{n}\right)$, $\alpha_{i}=\frac{\partial F}{\partial x_{i}}\left(x_{i}\right)$, and $f\left(x\right)=x-x_{i}$ (equivalent to $f(x)=x$). Therefore, we can use Eq. 8 and Theorem 5.3/5.4 to determine the point-wise error bounds from $F$. It is worth noticing that, in this case, for different $\left\{x_{i}\right\}$, we will have different $t$ from Theorem 5.3 and  5.4 as there are different $\alpha_{i}$ values from the partial derivatives.
In Algorithm 3, we combine the computation of point-wise data error bounds from multivariate QoI in either linear or non-linear format. When dealing with the linear format, the computation can often be simplified. For example, for linear multivariate QoIs, the $t$ in Line 14 can be shared over data regions.
@@ 5.4. QoI Error Validation and Correction
As previously described in Section 3.2, the design purpose of QPET is to optimize the compression ratio when strictly constraining the QoI error threshold. To fulfill this target, the point-wise and global error-bound calculations demonstrated in Section 5.1, Section 5.2, and Section 5.3 do not conservatively make the QoI value strictly preserved on each data element, but require a subsequent process to correct and store the out-of-bound data points (i.e., outliers). By selecting proper error-bound values, QPET can significantly reduce the average storage cost of each compressed data value, meanwhile introducing negligible overheads for correcting outliers as they are only a tiny portion ($<1\%$) of the whole input data. Specifically, QPET detects and corrects the outliers in the following scheme: Given the decompressed data (which are naturally acquired after the compression process), QPET computes the QoI values on it and compares them to the original QoI values. For out-of-bound univariate QoI errors, QPET losslessly encodes and stores all the corresponding original data values. For each out-of-bound multivariate QoI error, there are several related data values. To minimize the number of data values that are losslessly stored, QPET stores those values one by one and updates the QoI result after each value update until the QoI error gets bounded.
@@ 6.1.1. Experimental environment and datasets
We perform the evaluations on 6 real-world scientific datasets from diverse domains (details in Table 3). Experiments are operated on the Purdue Anvil computing cluster (Song et al., 2022) (each node is equipped with two 64-core AMD EPYC 7763 CPUs and 512GB DDR4 memory).
@@ 6.1.2. Baselines
Besides the QPET-integrated compressors, we included several existing solutions for QoI-preserving scientific lossy compression in the evaluations, which are in 2 categories: 1) Parameter-search-based solutions: general-purpose compressors (SZ3/HPEZ/SPERR) cannot directly bound specified QoI errors, so we apply parameter-search methods on top of them to figure out the best-fit (yielding highest compression ratio) data error bound for preserving the QoI. The parameter-search methods, including binary-search-based and FraZ (Underwood et al., 2020), are from OptZConfig (Underwood et al., 2022), the state-of-the-art scientific lossy compression parameter-search toolkit. When applying, we slightly revised those methods to get them better adapted and accelerated in the QoI-preserving tasks. Those baselines are named SZ3/HPEZ/SPERR-OptZ-R (R is short for revised), and they represent the best (fastest) results from both parameter-search methods on every base compressor. The parameter-search process often requires multiple iterations of data compression and QoI validation to fit the QoI error bound, making it typically quite slow. (2) Direct QoI-preserving solutions: (Jiao et al., 2022) provided the SZ3-based QoI-preserving compressor QoI-SZ3, and we further ported its QoI-preserving features to HPEZ, creating QoI-HPEZ. Moreover, we evaluated MGARD-QoI (Ainsworth et al., 2019). Those QoI-preserving compressors have limited support for diverse QoI formats. QoI-SZ3/HPEZ only supports square and logarithm (also their block averages), and MGARD-QoI only supports linear QoIs. So, we only tested them on the QoIs supported. Other existing QoI-integrated compressors are designed for different tasks (e.g., cpSZ (Liang et al., 2018a) only works for critical points in vector field data), so they are not included in our evaluation baselines.
@@ 6.1.3. QoI functions, experimental configurations, and evaluation metrics
Table 4 shows the QoI functions in the evaluation tasks. Among them, there are three different categories: point-wise, regional, and vector. They have diverse mathematical formats, and for many among them (such as $\tanh{x}$, $\frac{1}{n}\sum{x^{3}}$, and vector QoIs), QPET is the first framework that supports compression while preserving those QoIs. The selection of QoI functions in our evaluation is based on existing investigations and analysis (Jiao et al., 2022; Su et al., 2022; Wu et al., 2024) of QoIs in practical scientific data analysis tasks, such as physical transform (kinetic energy as velocity’s square), and clustering ($\sqrt{x^{2}+y^{2}+z^{2}}$ is the distance from origin when $x$, $y$, and $z$ are coordinates).
@@
For the OptZ-R parameter search, we set an early termination condition that triggers when a maximum QoI error between 90% and 100% of the required threshold is found, as it presents a near-optimal compression ratio with a reasonable search time. Regarding the compression configurations, we apply the default optimization level and compression-ratio-preferred mode for HPEZ. Regarding QPET parameters in Algorithms 3, we set $c=3$, $\beta=0.999$ for SPERR, $c=2$, $\beta=0.999$ for HPEZ, and $c=2$, $\beta=0.99999$ for SZ3. On SZ3 and HPEZ, the autocorrelation of decompression errors are relatively high when the error bound is large (Liu et al., 2022), so we linearly dynamically decrease $c$ when $\tau$ increases over $10^{-3}$, eventually to 1.0 when $\tau$ becomes $10^{-2}$.